This site is moving to  Please visit the new site and update any bookmarks you have. 

September 14, 2017 – Landsat 7 Anomaly Resolution

Engineers have successfully completed recovery efforts of the Landsat 7 Solid State Recorder (SSR) anomaly that occurred on September 13, 2017. The satellite resumed nominal imaging operation as of September 13 (DOY256) at 22:09 UTC (5:09 pm CST) and data are available to download from EarthExplorer, GloVis, or the LandsatLook Viewer. (Read More)

​September 13, 2017 - Landsat 7 Solid State Recorder Anomaly

Early morning September 13, 2017, the Landsat 7 Solid State Recorder (SSR) experienced an anomaly. The Flight Operations Team (FOT) initial analysis of telemetry indicated what is referred to as an "SSR choke anomaly" where the SSR locks up. This condition has occurred multiple times over the past several years. (Read more)

January 13, 2016 - Landsat 8 Thermal Data Update

Landsat 8 Thermal Infrared Sensor  (TIRS) data continue to be collected with the scene select mirror encoder electronics disabled (mode 0).  While in this mode, the TIRS line of sight model (LOS) will be regularly updated and modifications are being made to automate revisions to the LOS in the Landsat Product Generation System (LPGS). (Read More)

November 17, 2015 Landsat 8 TIRS Anomaly Update

The Landsat 8 Thermal Infrared Sensor (TIRS) Scene Select Mirror (SSM) encoder electronics anomaly investigation continues. In the meantime, Landsat 8 Level-1 products will continue to be made available consisting of valid data for the OLI solar reflective bands but with zero-fill data for the TIRS bands until corrections can be released in early 2016. (Read More)

November 3, 2015 - ​TIRS Scene Select Mirror Encoder Current Anomaly

At approximately 4:00 PM Central Standard Time (22:00 GMT) Sunday November 1, 2015 the TIRS instrument experienced an anomalous condition related to the instrument's ability to accurately measure the location of the Scene Select Mirror (SSM). (Read More)

October 28, 2015 - Landsat 8 Thermal Sensor Electronics Condition

The Landsat 8 Flight Operations Team is monitoring an increasing trend in the Thermal Infrared Sensor (TIRS) B-Side Encoder Electronics current. This condition is similar to the previous observations made on the Encoder A-Side electronics during the fall of 2014. (Read More)

October 13, 2015 - Recent Landsat 8 Data Acquired in Lower Truncation

In September and October 2015, some Landsat 8 scenes were acquired using the lower 12 bits of the entire 14-bit range. Scenes acquired in this mode support the study of dark features, but consequently cause bright data artifacts. For more information see the OLI Lower Truncation Webpage.

March 6, 2015 - Landsat 8 TIRS Sensor Resumes Nominal Operations

The Landsat 8 Thermal Infrared Sensor (TIRS) resumed normal imaging operations on March 4, 2015, and nominal blackbody and deep space calibration data collection will resume on March 7, 2015. (Read More)

February 26, 2015 - Landsat 8 Thermal Sensor Update

After extensive investigation and testing, the decision has been made to switch the TIRS Mechanism Control Electronics from the primary to redundant side on Monday, March 2, 2015. (Read More)

February 20, 2015 - Landsat 8 Thermal Sensor Update

The Landsat 8 satellite continues to acquire data from both onboard sensors, however Level-1 processing of Thermal Infrared Sensor (TIRS) data remains suspended – TIRS data bands currently consist of only zero-fill. (Read More)

January 27, 2015 - Landsat 8 TIRS Update

Landsat 8 continues to acquire OLI and TIRS data, although TIRS Level-1 processing remains suspended.

On December 19 the TIRS instrument on Landsat 8 was reconfigured due to detection of anomalous current levels associated with the scene select mirror encoder electronics. (Read More)

January 23, 2015 - Landsat 7 Data Update

The Landsat 7 Gap Phase Statistic (GPS) errors reported in the January 21, 2015 Headline have been corrected, and all data acquired on January 19, 2015 (DOY 19) forward have been reprocessed with valid GPS values.

January 21, 2015 - Landsat 8 Thermal Infrared (TIRS) data are still being collected but product generation remains suspended.

The Flight Operations Team and subject matter experts have conducted numerous tests in an effort to determine the cause of the current anomaly associated with the TIRS scan select mirror (SSM) encoder. (Read More)

January 21, 2015 - Recent Landsat 7 data changed

On Monday, January 19, 2015 (DOY 19) at 03:14:07 GMT, Gap Phase Statistic (GPS) errors were discovered in the compilation of Landsat 7 ETM+ Level-1 data products. Data processing was temporarily suspended until late afternoon January 20, 2015. (Read More)

December 22, 2014 - TIRS Level-1 Processing is Still Suspended

The configuration of the TIRS instrument on Landsat 8 remains unchanged since the detection of anomalous current levels associated with the scene select mirror (SSM) encoder electronics. (Read More)

December 19, 2014 - TIRS Calibration Reconfiguration

At 10:34 am (CST) the configuration of the TIRS instrument on Landsat 8 was reconfigured due to detection of anomalous current levels associated with the scene select mirror encoder electronics. (Read More)

November 4, 2014 - Landsat Ghostbusters

Shortly after the launch of Landsat 8, the Landsat calibration team noticed something strange: bright and dark stripes, or “banding” was showing up across certain images collected by the satellite’s Thermal Infrared Sensor (TIRS). NASA has provided a very informative article about these 'ghosting' artifacts.

September 29, 2014 - Landsat 7 Acquisitions Suspended

The Landsat 7 satellite stopped acquiring data Saturday afternoon, September 27, 2014 (DOY 270) due to an onboard telemetry unit issue. Nominal acquisitions are expected to resume late Monday, September 29, 2014 (DOY 272). More details will be provided when they become available.

March 15, 2011 - Short-term Suspension of Landsat 5 Acquisitions

On March 14, 2011, Landsat 5 experienced a minor anomaly. The spacecraft is currently operating normally, but it will take a few days to return the instrument to a nominal temperature. Data is expected to be available at the end of the week.

December 18, 2009 - Landsat 5 Anomaly

On Friday, December 18, 2009, the transmitter on Landsat 5 experienced technical difficulties. There has been no Landsat 5 data acquired since that time. The U.S. Geological Survey Flight Operations team has extensive experience with this transmitter and is currently addressing the situation.